• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name D. Parno

    Log entry time 12:19:42 on November 15, 2008

    Entry number 248993

    keyword=Scaler problem fixed (new Compton photon DAQ)

    We have fixed a problem with the Caen V560 scaler in the VME crate with the new Compton photon DAQ. The scaler had not been clearing properly (instead of going to zero, each channel went to 25890 when reset).

    We identified the problem as involving fixed bits. We moved the scaler to a different slot in the VME crate, making sure that it was firmly in place. The scaler channels now correctly reset to zero.

    A copy of this log entry has been emailed to: rom@jlab.org, gbfranklin@cmu.edu